OpenStudy (anonymous):

HELP ME PLEASE!!! The sum of two numbers is 54 . The larger number is 16 more than the smaller number. What are the numbers?

OpenStudy (anonymous):

i need the larger number and the smaller number

OpenStudy (anonymous):

@SolomonZelman help me plzzzzz

OpenStudy (anonymous):

dgdfggdggdg

OpenStudy (anonymous):

drgdrg

OpenStudy (anonymous):

g

OpenStudy (anonymous):

d

OpenStudy (anonymous):

rgd

OpenStudy (anonymous):

dggdr

OpenStudy (anonymous):

@dtan5457 x=36

OpenStudy (solomonzelman):

`The sum of two numbers is 54 .` \(\large\color{black}{x+y=54 }\) `The larger number is 16 more than the smaller number.` \(\large\color{black}{y=x+16 }\)

OpenStudy (dtan5457):

The bigger number is 36, what is the smaller one then?

OpenStudy (solomonzelman):

dtan, where did you get 56, and 56-x from, may I know?

OpenStudy (anonymous):

smaller number will be 18?

OpenStudy (anonymous):

@SolomonZelman is this wrong??

OpenStudy (dtan5457):

Time out. Read the question wrong, it should be 54-x

OpenStudy (solomonzelman):

apparently I saved you, blueBab, :P

OpenStudy (dtan5457):

Let x=the bigger number Let (54-x)=smaller number (54-x)=x-16 NOW, solve for x..lol

OpenStudy (anonymous):

@SolomonZelman thanks ur my hero =D

OpenStudy (solomonzelman):

not a hero, but... anyways, can you proceed? \(\large\color{black}{y+x=54 }\) \(\large\color{black}{y=x+16 }\) \(\large\color{black}{x+16+x=54 }\) \(\large\color{black}{2x+16=54 }\) \(\large\color{black}{x=? }\)

OpenStudy (camerondoherty):

Look, heres to make it easier you have 54 subtract 16 divide by 2 then add 16 to one of them and that one is the larger, the other is the smaller c:

OpenStudy (anonymous):

x=19 @SolomonZelman